pleasee use sign chart
(5x-2)(x+4)2 x-5 0.
Answer #1

152-2 out 09 LO x-5 10 11 ² (5x-2)(x+4)² = 0 52-2=0 E: (x+4)2710 12 x = 2/5 Ans. 01